OpenStudy (anonymous):

in a polynomial with degree 3, can we find a simple expression for the sum of the cubes of the three roots?

OpenStudy (anonymous):

In polynomial ax^3+bx^2+cx+d = 0 3 roots are p, q and r p+q+r = -b/a pq+pr+qr = c/a pqr = -d/a can we find a simple expression for p^3+q^3+r^3 using the above identities?

OpenStudy (anonymous):

For example, in polynomial ax^2+bx+c = 0 roots are p and q p+q = -b/a pq = c/a p^2+q^2 = (p+q)^2 - 2pq

OpenStudy (inkyvoyd):

eh?

OpenStudy (inkyvoyd):

I doubt it. Do you know how to solvethe cubic?

OpenStudy (anonymous):

what do you mean by solve the cubic?

OpenStudy (asnaseer):

\[\begin{align} (p+q+r)^3&=p^3+q^3+r^3+3p^2q+3p^2r+3pq^2+6pqr+3pr^2+3q^2r+3qr^2\\ &=p^3+q^3+r^3+3(p^2q+p^2r+pq^2+2pqr+pr^2+q^2r+qr^2)\\ &=p^3+q^3+r^3+3(p(pq+pr+qr)+q(pq+pr+qr)+pr^2+qr^2)\\ &=p^3+q^3+r^3+3((p+q)(pq+pr+qr)+pr^2+qr^2)\\ &=p^3+q^3+r^3+3((p+q+r)(pq+pr+qr)-r(pq+pr+qr)+pr^2+qr^2)\\ &=p^3+q^3+r^3+3((p+q+r)(pq+pr+qr)-pqr) \end{align}\]therefore:\[p^3+q^3+r^3=(p+q+r)^3-3((p+q+r)(pq+pr+qr)-pqr)\] I think that is correct.

OpenStudy (inkyvoyd):

Solve the cubic equation for its 3 roots.

OpenStudy (anonymous):

@inkyvoyd, that's what we're trying to avoid

OpenStudy (anonymous):

Thanks heaps, this is now closed

OpenStudy (asnaseer):

yw

OpenStudy (inkyvoyd):

@psujono , medal asnaseer if you think answer was good!

OpenStudy (asnaseer):

thx for the thought inkyvoyd, but I use this site to learn and teach rather than gain any medals, so I am not really bothered if I don't get a medal. I just enjoy the interaction with other users. :)

OpenStudy (anonymous):

I'll medal you anyway, good answer - thanks!

OpenStudy (asnaseer):

thx

OpenStudy (inkyvoyd):

@asnaseer , there are people like me that prefer medals (LOL), which is why I reminded him to medal (so, you don't forget xD, because if someone like me helps you out, the will rage without a medal)

OpenStudy (asnaseer):

@inkyvoyd - Good point - I guess I should think of everyone else using this site as well :)

OpenStudy (inkyvoyd):

No, just me LOL